### v1

@@ Abstract
Index Terms:
Artificial Intelligence Safety, Autonomous agents, Barrier Functions, Intelligent control, Lyapunov Functions, Neural Networks (NNs), Reinforcement Learning.
@@ I Introduction
Control Lyapunov functions (CLF) [37] and control barrier functions (CBF) [38] have proven to be an effective way to guarantee safety through stability of a closed-loop system and defining safe sets, respectively. Ensuring safe operation for dynamical systems involving human interaction, expensive equipment etc., requires strong guarantees during and after the learning process for RL agents. Using CBFs and CLFs to provide stability and safety guarantees is especially well suited for learning because both features may be stated using learnable functions. Considering their usefulness in a learning framework and their applicability to a wide variety of control frameworks, this paper reviews methods using Lyapunov and barrier functions to ensure safety and stability guarantees of a computed policy while using RL for decision-making problems. Furthermore, most real-world applications require safety during the training process (for example, navigation in unknown environments for drones, human presence in application etc.) which makes this problem crucial as it provides theoretical guarantees for safety. The key motivation for this review is to discuss methods employed in recent work, not covered in reviews and surveys mentioned in Section II and particularly focus on use of Lyapunov and barrier functions for RL to guarantee satisfying system constraints and stability in terms of the computed policy. We focus more on the theoretical approaches present in literature rather than the type of applications it is employed. To the best of author’s knowledge, such a review focusing on use of Lyapunov and barrier functions in safe RL does not exist and this research gap is addressed in this review. Rest of the review is organized as follows: Section II presents scope of work for different reviews and surveys considered. Section III discusses preliminary theory and notations briefly to provide the reader a background for further discussion. Section IV presents and summarizes the current and prior work related to use of: (i) Lyapunov functions in RL; (ii) Barrier functions in RL; (iii) Lyapunov and barrier functions in RL. Section V provides a discussion on future research directions, current challenges and potential real-world applications. Finally, Section VI presents conclusions to the review conducted.
II Related Work
@@ III-A Definition of Safety
A broad definition of safety is taken into account for a dynamical system that is being regulated by a control policy. In order for the system to be considered safe, it must be ensured that the system state trajectories will never enter any unsafe set under the existing control policy. By guaranteeing a safe set’s forward invariance, safety can be ensured. That is, for every any time $t$ ($t\geq 0$), all trajectories beginning in the set of safe states will remain there. Formally, given an unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ that does not contain the set of initial conditions $\mathcal{X}_{0}$. Safety is then defined for the system if it never enters the unsafe set $\mathcal{X}_{u}$ when starting within $\mathcal{X}_{0}$. The notion of safety can be further extended in terms of stability in the Lyapunov sense, which will be discussed in Section III-C. Fig. 1 provides a notion of safe set.
@@
III-B Constrained Markov Decision Process (CMDP) and Safe RL
@@ III-C Lyapunov Functions
where $x_{d}\in\mathcal{X}$ is the desired state for the system and $f_{cl}$ is the closed loop dynamics of the system. $\nabla V(x)f_{cl}(x)$ denotes the Lie derivative of $V$ along the function $f_{cl}$. Control Lyapunov functions provide formal guarantees for stabilizability of open-loop systems, which implies the existence of a controller that drives the closed loop system to stability. Formally, a control Lyapunov function (CLF) is a criterion that provides asymptotically stabilizing controllers for a general nonlinear dynamical system, which was first generalized by Sontag [48]. For a control affine form in Section III-A, a CLF $V$ is a smooth (continuous and differentiable up to some order), proper ($V(x)\rightarrow\infty$ as $\|x\|\rightarrow\infty$) and positive definite ($V(x)>0$ and $V(0)=0$ for $x\neq 0$) function:
@@ III-D Barrier Functions
IV Literature Review
IV-A Lyapunov Functions for Reinforcement Learning
The closed-loop system’s stability assurances can be used to ensure safety guarantees. These methods frequently rely on control Lyapunov functions (CLFs) for Lyapunov stability verification presented in Section III-C. CLFs were formalized for a general nonlinear dynamical system by Sontag [48]. An interesting property that a lot of work in safe RL leverages is the fact that level sets of CLFs are both attractive and invariant. Fig. 4 shows a brief classification of using Lyapunov functions approaches in RL. The earliest work using Lyapunov functions with RL was proposed by Perkins et al. [51], where the authors use Lyapunov functions to design multiple controllers which the control policy can switch between, which is designed using RL (agent decides when to switch between Lyapunov-based control laws). The system is formulated as a MDP and the authors implement state-action-reward-state-action (SARSA) ($\lambda$) RL algorithm with cerebellar model articulation controller (CMAC) function approximator [52] for action-value functions. The use of Lyapunov function ensures safety guarantees for any switching policy and furthermore, empirically showed accelerated learning.
@@ IV-A Lyapunov Functions for Reinforcement Learning
TABLE I: Summary of safe RL approaches with Lyapunov functions
@@
Stochastic gradient descent is used to minimize the loss function. The proposed algorithm, “Policy optimization with self-learned almost Lyapunov critics algorithm”, follows a structure similar to PPO with the addition of Lyapunov risk minimization for the parameter $\theta$ updates. The algorithm’s performance is compared with Lyapunov actor-critic, soft actor critic and PPO for inverted pendulum, quadrotor control, automobile path-tracking and MuJoCo Hopper. A similar approach is proposed in [63, 64]. Westenbroek et al. [72] propose use of a CLF for an inaccurate model of the system and use the rate of dissipation of the CLF as a constraint (threshold for minimum acceptable rate) on closed-loop controller. The controller parameters are optimized using model-free policy optimization using data sampled from the system. The framework is tested on a double pendulum and underactuated bipedal robots using soft actor-critic (SAC) algorithm. Westenbroek et al. [69] also propose use of CLF as a value function and show that smaller values for discount factor lead to stabilizing controllers with the choice of CLF as the value function. The proposed algorithm is tested on cartpole and an A1 quadruped using the SAC algorithm with augment value function approach. Authors in [65], train a DDPG agent based on the violations of Lyapunov conditions for nonlinear vehicle dynamics model in a lane-following environment. The two-step approach first learns system dynamics and a CLF using NNs for a current policy and secondly, trains a DDPG agent to maximize reward and minimize violation of Lyapunov conditions using the learned CLF.
@@
Table I summarizes the literature reviewed for use of Lyapunov functions in RL. It can be observed that two formal approaches to incorporate Lyapunov functions with RL involve: (i) Lyapunov-based reward-shaping; (ii) CLF-RL-QP formulations; (iii) candidate Lyapunov function as the value/critic function; (iv) Lyapunov/CLF-based loss function. The key challenge lies in approximating or finding a suitable candidate Lyapunov function without having prior or imperfect information of the model/environment. Using model-based methods can circumvent this problem by providing a candidate Lyapunov function based on system dynamics, but its suitability for model-free approaches remains an open area of research. Another difference in methodology for these approaches depend on approximating the Lie derivative of Lyapunov function along the system dynamics, which requires some prior knowledge or is approximated in a data-driven format. Lie derivatives are critical in computing the gradient of loss function with Lyapunov/CLF terms.
IV-B Barrier Functions for Reinforcement Learning
TABLE II: Summary of safe RL approaches with Barrier functions
@@ IV-B Barrier Functions for Reinforcement Learning
There has been significant recent research on use of barrier functions with RL to define safe sets for the control policy. Fig. 5 shows a broad classification for the use of barrier function-based approaches in safe RL. Similar to reward-shaping using CLF/Lyapunov terms in Section IV-A, reward-shaping using CBF terms is observed in several works to ensure state constraints are satisfied. Marvi et al. [91] propose adding a CBF term to the cost-to-go function for a safe optimal control problem with nonlinear dynamics. A coefficient $\gamma$ (damping term) is introduced along with the CBF term to control the effect of barrier term, this can be interpreted as a trade-off between safety and optimality. The authors also prove the additional barrier term maintains safety and optimality (under the assumption of finite value barrier function and suitable value of $\gamma$). The MPC formulation for a discrete time control system is augmented as follows:
@@
In this case, dependence on all prior CBF controller enhances learning efficiency. The second approach guides exploration as well by taking prior computed CBFs into consideration. Probabilistic guarantees are provided for TRPO and DDPG algorithms implemented. In [97], authors propose a generalized CBF for a model-based constrained RL policy optimization algorithm. The problem is formulated for an actor-critic algorithm with the actor policy update as an optimization problem with CBF constraints. Lagrange multipliers are used to reformulate the problem as an unconstrained optimization problem along with a distance constraint (collision avoidance simulations). Furthermore, a parameter $\xi$ (constraint violation metric) is computed to compare with a predetermined threshold for updating the parameters of CBF to adjust constraint satisfaction. A robust CBF is proposed in [96] for a nonlinear system with disturbance. The disturbance is approximated using a GP (this approximated model is used to compute the Lie derivatives in (12)) and the robust CBF is combined with RL actions to form a relaxed QP problem. The slack variable is introduced to avoid deadlock when no safe action is available. The proposed approach is tested on unicycle and car following environments using a SAC-Robust CBF formulation. Similar optimization formulations are observed in [102, 103, 104, 99, 100, 106]. Soft-barrier constraints are proposed in [101], where a slack variable $\gamma$ is introduced to relax the barrier constraints and ensure a feasible solution for the CBF-RL-QP formulation. A multi-agent RL with decentralized CBF shields is proposed by authors in [98] for a collision-avoidance problem. The two types decentralized CBF shields are formulated namely, cooperative and non-cooperative, the safe policy is computed as a constrained optimization problem for the two types of CBF shields and action by the RL agent, similar to (20). A multi-agent DDPG-CBF formulation is tested on a two patrolman task. In [104], authors use disturbance observers and CBFs with RL to provide safety guarantees. A policy is sampled from the RL agent followed by disturbance estimation from the disturbance-based observer, similar to prior approaches discussed Lie derivative is approximated using disturbance-based observers. A QP (20) is solved using the CBF as constraints and the action sampled by the RL agent to solve for the safe action. SAC algorithm is used for simulations on a 2D quadrotor and a unicycle model. Hou et al. [105] and Kalaria et al. [106] follow a similar disturbance-observer-based CBF filter enforce hard constraints.
@@
It can be seen that most approaches concerning use of barrier functions follow a relaxed QP formulation to act as a compensating controller. The key differences in approaches are how the nonlinear dynamics are approximated to compute the safe set (forward invariant set). System information becomes necessary to compute these safe sets and conservative policies due to restricted exploration is observed. Table II summarizes the literature reviewed for use of barrier functions in safe RL.
@@
IV-C Lyapunov and Barrier Functions for Reinforcement Learning
@@
TABLE III: Summary of safe RL approaches with Lyapunov and Barrier functions
@@
Table III summarizes the literature reviewed for cumulative use of Lyapunov and barrier functions in RL. Most methods combining Lyapunov and barrier functions employ two of the following methods:(i) Solving a constrained optimization problem with Lyapunov and barrier constraints; (ii) using Lyapunov-barrier terms in the critic/actor loss function; (iii) Lyapunov-barrier-based reward shaping. It remains an open question for the feasibility and complexity of solving a constrained optimization problem at each iteration whilst optimizing the parameters for RL in case of large-scale problems, to satisfy the cumulative Lyapunov-Barrier constraints. A similar trend can also be observed in terms of selection for a candidate Lyapunov function for model-free approaches. Proving the effectiveness of approximating a candidate Lyapunov function with sampled data from the environment remains an open area of research.
V Discussion
Safe RL with Lyapunov or barrier functions is an effective and promising field that aims to ensure the safety of an agent while it learns a task through RL. Table III and III summarize the methodology of using Lyapunov and barrier functions in RL.
V-A Simulation Benchmarks
@@ V-D Future Work
V-E Potential and Current Applications

### v2

@@ Abstract
{IEEEImpStatement}
Reinforcement learning and data-driven control has seen significant research over the last two decades given its usability and effectiveness for computing control policies in unknown stochastic environments. Given most real-world scenarios are difficult to model and further have significant unknown disturbances, the problem becomes especially challenging. Safe reinforcement learning provides an elegant solution to these challenges as agents can safely explore and tackle these environments which creates several use cases for real-world applications as covered in this review.
Reinforcement learning in its bare-bones fails to satisfy or provide formal guarantees for safety and stability with respect to the computed policy. Lyapunov and barrier functions have long been employed in control theory and optimization problems to guarantee system stability and satisfy state constraints, respectively. Incorporating Lyapunov and barrier methods in reinforcement learning helps to address the issues of safety and stability to increase and promote its use in safety critical real-world applications.
The advantages of safe reinforcement learning extend directly to real-world applications and provide exciting opportunities to deploy agents which can learn safely in real-time, as discussed and covered in this review. Most of the literature in this field is fairly recent, dating back to 2017. This review compiles and presents important theoretical developments and challenges in the field and, further proposes possible directions for improving existing formulations.
{IEEEkeywords}
Reinforcement Learning, Lyapunov Functions, Barrier Functions, Neural Networks (NNs), Artificial Intelligence Safety, Autonomous agents, Intelligent control.
1 Introduction
@@
Control Lyapunov functions (CLF) [37] and control barrier functions (CBF) [38] have proven to be an effective way to guarantee safety through stability of a closed-loop system and defining safe sets, respectively. Ensuring safe operation for dynamical systems involving human interaction, expensive equipment etc., requires strong guarantees during and after the learning process for RL agents. Using CBFs and CLFs to provide stability and safety guarantees is especially well suited for learning because both features may be stated using learnable functions. Considering their usefulness in a learning framework and their applicability to a wide variety of control frameworks, this paper reviews methods using Lyapunov and barrier functions to ensure safety and stability guarantees of a computed policy while using RL for decision-making problems. Furthermore, most real-world applications require safety during the training process (for example, navigation in unknown environments for drones, human presence in application etc.) which makes this problem crucial as it provides theoretical guarantees for safety. The key motivation for this review is to discuss methods employed in recent work, not covered in reviews and surveys mentioned in Section 2 and particularly focus on use of Lyapunov and barrier functions for RL to guarantee satisfying system constraints and stability in terms of the computed policy. We focus more on the theoretical approaches present in literature rather than the type of applications it is employed. To the best of author’s knowledge, such a review focusing on use of Lyapunov and barrier functions in safe RL does not exist and this research gap is addressed in this review. Rest of the review is organized as follows: Section 2 presents scope of work for different reviews and surveys considered. Section 3 discusses preliminary theory and notations briefly to provide the reader a background for further discussion. Section 4 presents and summarizes the current and prior work related to use of: (i) Lyapunov functions in RL; (ii) Barrier functions in RL; (iii) Lyapunov and barrier functions in RL. Section 5 provides a discussion on future research directions, current challenges and potential real-world applications. Finally, Section 6 presents conclusions to the review conducted.
2 Related Work
@@ 3.1 Definition of Safety
A broad definition of safety is taken into account for a dynamical system that is being regulated by a control policy. In order for the system to be considered safe, it must be ensured that the system state trajectories will never enter any unsafe set under the existing control policy. By guaranteeing a safe set’s forward invariance, safety can be ensured. That is, for every any time $t$ ($t\geq 0$), all trajectories beginning in the set of safe states will remain there. Formally, given an unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ that does not contain the set of initial conditions $\mathcal{X}_{0}$. Safety is then defined for the system if it never enters the unsafe set $\mathcal{X}_{u}$ when starting within $\mathcal{X}_{0}$. The notion of safety can be further extended in terms of stability in the Lyapunov sense, which will be discussed in Section 3.3. Fig. 1 provides a notion of safe set.
@@
3.2 Constrained Markov Decision Process (CMDP) and Safe RL
@@ 3.3 Lyapunov Functions
where $x_{d}\in\mathcal{X}$ is the desired state for the system and $f_{cl}$ is the closed loop dynamics of the system. $\nabla V(x)f_{cl}(x)$ denotes the Lie derivative of $V$ along the function $f_{cl}$. Control Lyapunov functions provide formal guarantees for stabilizability of open-loop systems, which implies the existence of a controller that drives the closed loop system to stability. Formally, a control Lyapunov function (CLF) is a criterion that provides asymptotically stabilizing controllers for a general nonlinear dynamical system, which was first generalized by Sontag [48]. For a control affine form in Section 3.1, a CLF $V$ is a smooth (continuous and differentiable up to some order), proper ($V(x)\rightarrow\infty$ as $\|x\|\rightarrow\infty$) and positive definite ($V(x)>0$ and $V(0)=0$ for $x\neq 0$) function:
@@ 3.4 Barrier Functions
4 Literature Review
4.1 Lyapunov Functions for Reinforcement Learning
The closed-loop system’s stability assurances can be used to ensure safety guarantees. These methods frequently rely on control Lyapunov functions (CLFs) for Lyapunov stability verification presented in Section 3.3. CLFs were formalized for a general nonlinear dynamical system by Sontag [48]. An interesting property that a lot of work in safe RL leverages is the fact that level sets of CLFs are both attractive and invariant. Fig. 4 shows a brief classification of using Lyapunov functions approaches in RL. The earliest work using Lyapunov functions with RL was proposed by Perkins et al. [51], where the authors use Lyapunov functions to design multiple controllers which the control policy can switch between, which is designed using RL (agent decides when to switch between Lyapunov-based control laws). The system is formulated as a MDP and the authors implement state-action-reward-state-action (SARSA) ($\lambda$) RL algorithm with cerebellar model articulation controller (CMAC) function approximator [52] for action-value functions. The use of Lyapunov function ensures safety guarantees for any switching policy and furthermore, empirically showed accelerated learning.
@@ 4.1 Lyapunov Functions for Reinforcement Learning
Table 1: Summary of safe RL approaches with Lyapunov functions
@@
Stochastic gradient descent is used to minimize the loss function. The proposed algorithm, “Policy optimization with self-learned almost Lyapunov critics algorithm”, follows a structure similar to PPO with the addition of Lyapunov risk minimization for the parameter $\theta$ updates. The algorithm’s performance is compared with Lyapunov actor-critic, soft actor critic and PPO for inverted pendulum, quadrotor control, automobile path-tracking and MuJoCo Hopper. A similar approach is proposed in [61, 62]. Westenbroek et al. [71] propose use of a CLF for an inaccurate model of the system and use the rate of dissipation of the CLF as a constraint (threshold for minimum acceptable rate) on closed-loop controller. The controller parameters are optimized using model-free policy optimization using data sampled from the system. The framework is tested on a double pendulum and under-actuated bipedal robots using soft actor-critic (SAC) algorithm. Westenbroek et al. [68] also propose use of CLF as a value function and show that smaller values for discount factor lead to stabilizing controllers with the choice of CLF as the value function. The proposed algorithm is tested on cartpole and an A1 quadruped using the SAC algorithm with augment value function approach. Authors in [63], train a DDPG agent based on the violations of Lyapunov conditions for nonlinear vehicle dynamics model in a lane-following environment. The two-step approach first learns system dynamics and a CLF using NNs for a current policy and secondly, trains a DDPG agent to maximize reward and minimize violation of Lyapunov conditions using the learned CLF.
@@
Table 1 summarizes the literature reviewed for use of Lyapunov functions in RL. It can be observed that two formal approaches to incorporate Lyapunov functions with RL involve: (i) Lyapunov-based reward-shaping; (ii) CLF-RL-QP formulations; (iii) candidate Lyapunov function as the value/critic function; (iv) Lyapunov/CLF-based loss function. The key challenge lies in approximating or finding a suitable candidate Lyapunov function without having prior or imperfect information of the model/environment. Using model-based methods can circumvent this problem by providing a candidate Lyapunov function based on system dynamics, but its suitability for model-free approaches remains an open area of research. Another difference in methodology for these approaches depend on approximating the Lie derivative of Lyapunov function along the system dynamics, which requires some prior knowledge or is approximated in a data-driven format. Lie derivatives are critical in computing the gradient of loss function with Lyapunov/CLF terms.
4.2 Barrier Functions for Reinforcement Learning
Table 2: Summary of safe RL approaches with Barrier functions
@@ 4.2 Barrier Functions for Reinforcement Learning
There has been significant recent research on use of barrier functions with RL to define safe sets for the control policy. Fig. 5 shows a broad classification for the use of barrier function-based approaches in safe RL. Similar to reward-shaping using CLF/Lyapunov terms in Section 4.1, reward-shaping using CBF terms is observed in several works to ensure state constraints are satisfied. Marvi et al. [91] propose adding a CBF term to the cost-to-go function for a safe optimal control problem with nonlinear dynamics. A coefficient $\gamma$ (damping term) is introduced along with the CBF term to control the effect of barrier term, this can be interpreted as a trade-off between safety and optimality. The authors also prove the additional barrier term maintains safety and optimality (under the assumption of finite value barrier function and suitable value of $\gamma$). The MPC formulation for a discrete time control system is augmented as follows:
@@
In this case, dependence on all prior CBF controller enhances learning efficiency. The second approach guides exploration as well by taking prior computed CBFs into consideration. Probabilistic guarantees are provided for TRPO and DDPG algorithms implemented. In [97], authors propose a generalized CBF for a model-based constrained RL policy optimization algorithm. The problem is formulated for an actor-critic algorithm with the actor policy update as an optimization problem with CBF constraints. Lagrange multipliers are used to reformulate the problem as an unconstrained optimization problem along with a distance constraint (collision avoidance simulations). Furthermore, a parameter $\xi$ (constraint violation metric) is computed to compare with a predetermined threshold for updating the parameters of CBF to adjust constraint satisfaction. A robust CBF is proposed in [96] for a nonlinear system with disturbance. The disturbance is approximated using a GP (this approximated model is used to compute the Lie derivatives in (12) and the robust CBF is combined with RL actions to form a relaxed QP problem. The slack variable is introduced to avoid deadlock when no safe action is available. The proposed approach is tested on unicycle and car following environments using a SAC-Robust CBF formulation. Similar optimization formulations are observed in [102, 103, 104, 99, 100, 106]. Soft-barrier constraints are proposed in [101], where a slack variable $\gamma$ is introduced to relax the barrier constraints and ensure a feasible solution for the CBF-RL-QP formulation. A multi-agent RL with decentralized CBF shields is proposed by authors in [98] for a collision-avoidance problem. The two types decentralized CBF shields are formulated namely, cooperative and non-cooperative, the safe policy is computed as a constrained optimization problem for the two types of CBF shields and action by the RL agent, similar to (20). A multi-agent DDPG-CBF formulation is tested on a two patrolman task. In [104], authors use disturbance observers and CBFs with RL to provide safety guarantees. A policy is sampled from the RL agent followed by disturbance estimation from the disturbance-based observer, similar to prior approaches discussed Lie derivative is approximated using disturbance-based observers. A QP (20) is solved using the CBF as constraints and the action sampled by the RL agent to solve for the safe action. SAC algorithm is used for simulations on a 2D quadrotor and a unicycle model. Hou et al. [105] and Kalaria et al. [106] follow a similar disturbance-observer-based CBF filter enforce hard constraints.
@@
It can be seen that most approaches concerning use of barrier functions follow a relaxed QP formulation to act as a compensating controller. The key differences in approaches are how the nonlinear dynamics are approximated to compute the safe set (forward invariant set). System information becomes necessary to compute these safe sets and conservative policies due to restricted exploration is observed. Table 2 summarizes the literature reviewed for use of barrier functions in safe RL.
@@
4.3 Lyapunov and Barrier Functions for Reinforcement Learning
@@
Table 3: Summary of safe RL approaches with Lyapunov and Barrier functions
@@
Table 3 summarizes the literature reviewed for cumulative use of Lyapunov and barrier functions in RL. Most methods combining Lyapunov and barrier functions employ two of the following methods:(i) Solving a constrained optimization problem with Lyapunov and barrier constraints; (ii) using Lyapunov-barrier terms in the critic/actor loss function; (iii) Lyapunov-barrier-based reward shaping. It remains an open question for the feasibility and complexity of solving a constrained optimization problem at each iteration whilst optimizing the parameters for RL in case of large-scale problems, to satisfy the cumulative Lyapunov-Barrier constraints. A similar trend can also be observed in terms of selection for a candidate Lyapunov function for model-free approaches. Proving the effectiveness of approximating a candidate Lyapunov function with sampled data from the environment remains an open area of research.
5 Discussion
Safe RL with Lyapunov or barrier functions is an effective and promising field that aims to ensure the safety of an agent while it learns a task through RL. Table 12 and 3 summarize the methodology of using Lyapunov and barrier functions in RL.
5.1 Simulation Benchmarks
@@ 5.4 Future Work
5.5 Potential and Current Applications